OpenStudy (anonymous):

Translate into equation: The angles of a triangle if one angle is one-fourth the second, and the third is the sum of the first two.

OpenStudy (anonymous):

the total is \(180\)

OpenStudy (anonymous):

if you call one \(x\) then the other is \(\frac{1}{4}x\) and the third is \(x+\frac{1}{4}x\) add them and get \[x+\frac{1}{4}x+x+\frac{1}{4}x=180\\ 2\tfrac{1}{2}x=180\]